OpenStudy (anonymous):

square root of 100v^3

OpenStudy (mertsj):

\[\sqrt{100v^3}=\sqrt{100}\sqrt{v^2v}=10v \sqrt{v}\]

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw